OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -9) and a directrix y = 9. answer choices: y = -1/9^2 y^2 = -36x y = -1/36^2 y^2 = -9x

OpenStudy (jdoe0001):

where do you think is the origin at ?

OpenStudy (anonymous):

Is the origin also the vertex?

OpenStudy (jdoe0001):

 so yes, is at the origin, notice that you can just plug those values in a "focus form" of a parabola like, the parabola is going downwards, so the \(\large \bf (x-h)^2=-4p(y-k)\) (h, k) = vertex coordinates p = distance from the vertex to the focus or directrix, since the directrix and focus are equidistant from the vertex

OpenStudy (anonymous):

Okay so, I would no be able to fill in the values. (x-0)^2=-4(9)(y-0) x^2=-36y

OpenStudy (jdoe0001):

\(\bf x^2=-36y\implies -\cfrac{x^2}{36}=y\) yeap

OpenStudy (anonymous):

Awesome! Thank you so much for helping! ~clear skies

OpenStudy (jdoe0001):

yw